OpenStudy (anonymous):

air contains 78% N2, 21% O2, and 1% Ar, by volume. what is the density of air at 1,000. torr and -10C?

OpenStudy (mos1635):

PV=nRT PV=(m/MW)RT P(V/m)=(R/MW)T P(1/ρ)=(R/Mw)T Mw = is the wean value of gas = air Mw=0.78*Ar(N) + 0.28 Ar(O) + 0.01 Ar (Ar) Ar=atomic relative mass

OpenStudy (mos1635):